# Federated Learning over Connected Modes

### v2

@@ Abstract
Statistical heterogeneity in federated learning poses two major challenges: slow global training due to conflicting gradient signals, and the need of personalization for local distributions. In this work, we tackle both challenges by leveraging recent advances in linear mode connectivity — identifying a linearly connected low-loss region in the weight space of neural networks, which we call solution simplex. We propose federated learning over connected modes (Floco), where clients are assigned local subregions in this simplex based on their gradient signals, and together learn the shared global solution simplex. This allows personalization of the client models to fit their local distributions within the degrees of freedom in the solution simplex and homogenizes the update signals for the global simplex training. Our experiments show that Floco accelerates the global training process, and significantly improves the local accuracy with minimal computational overhead.
@@ 1 Introduction
Federated learning (FL) [1] is a decentralized machine learning paradigm that facilitates collaborative model training across distributed devices while preserving data privacy. However, in typical real applications, statistical heterogeneity—non-identically and independently distributed (non-IID) data distributions at clients—makes it difficult to train well-performing models. To tackle this difficulty, various methods have been proposed, e.g., personalized FL [2], clustered FL [3], advanced client selection strategies [4], robust aggregation [5], and federated meta- and multi-task learning approaches [6]. These methods aim either at training a global model that performs well on the global distribution [7], or, as it is common in personalized FL, at training multiple client-dependent models each of which performs well on its local distribution [8]. These two aims often pose a trade-off—a model that shows better global performance tends to show worse local performance, and vice versa. In this work, we aim to develop a FL method that improves both global and local performance.
@@
Our approach is leverages recent findings on mode connectivity [9, 10, 11]—the existence of low-loss paths in the weight parameter space between independently trained neural networks—and its applications [12]. These works show that minima for the same task are typically connected by simple low-loss curves, and that such connectivity benefits training for multi-task and continual learning. More specifically, the authors show that embracing such mode connectivity between models improves accuracy on each task and remedies the risk of catastrophic forgetting.
@@
Figure 1: Floco expresses each client as a point ($\star$ in the top-middle plot) by projecting the gradient signals onto the simplex, so that similar clients are close to each other. In each communication round, each client uniformly samples points in the neighborhood of their projected point (top-right plot), and jointly train the solution simplex. The lower row plots the resulting test loss on the solution simplex, where the loss for the global distribution (left) is uniformly small, while the losses for individual local distributions (middle for client 1 and right for client 2) are small around their projected points.
@@
In this paper, we leverage such effects, and propose federated learning over connected modes (Floco), where the clients share and together train a solution simplex—a lineraly connected low-loss region in the weight parameter space. Specifically, Floco represents clients as points within the standard simplex based on the similarity between their gradients, and assigns each client a specific subregion of the simplex. Clients then participate in FL by sampling different models within their assigned subregions and sending back the gradient information to update the vertices of the global solution simplex (see Fig.1). This method facilitates collaborative training through the common solution simplex, while allowing for client-specific personalization according to their local data distributions.
@@
Our experiments show that Floco outperforms state-of-the-art approaches in both global (FedProx [13]), and personalized (Ditto [14]) FL without introducing significant computational overhead. We also show other benefits of Floco, including better uncertainty estimation, improved worst client performance, and smaller divergence of the gradient signals.
@@
  • We propose Floco, a novel FL method that trains a solution simplex for mitigating the statistical heterogeneity of clients, and demonstrate its state-of-the-art performance for global and local personalized FL.
  • We propose a simple projection method to express clients as points in the standard simplex based on the gradient signals, and establish a procedure of subregion assignements.
  • We conduct experimental evaluations on semi-artificial and real-world FL benchmarks with detailed analyses of the behavior of Floco, which give insights into the mechanism on how it improves performance compared to the baselines.
Our code is submitted as a supplement and will be made public upon acceptance.
@@ 2.1 Federated Learning
Here, $p(k)$ is the normalized population of data samples over the clients, $p_{k}(\boldsymbol{x},y)$ is the data distribution for the client $k$, and $f(g,(\boldsymbol{x},y))$ is the loss, e.g., cross-entropy, of the model $g$ on a sample $(\boldsymbol{x},y)\in\mathbb{R}^{I}\times\{1,\ldots,L\}$, where $I$ is the dimension of an input data sample. Global [15] and personalized [8] FL aim to approximate $g_{0}^{*}$ and $\{g_{k}^{*}\}_{k=1}^{K}$, respectively, by using the training data $\mathcal{D}=\{\mathcal{D}_{k}\}_{k=1}^{K}$ observed by the clients. Throughout the paper, we assume that all models are neural networks (NNs) $\widehat{y}=g_{k}(\boldsymbol{x};\boldsymbol{w}_{k})$ with the same architecture, and represent the model $g_{k}$ with its NN weight parameters $\boldsymbol{w}_{k}\in\mathbb{R}^{D}$, i.e., we hereafter represent $g_{k}(\boldsymbol{x};\boldsymbol{w}_{k})$ by $\boldsymbol{w}_{k}$ and thus denote, e.g., $F_{k}^{*}(g_{k})$ by $F_{k}^{*}(\boldsymbol{w}_{k})$. Let $N=\sum_{k=1}^{K}N_{k}$ be the total number of samples, where $N_{k}=|\mathcal{D}_{k}|$.
@@
where $\lambda$ controls the divergence from the global model.
Ditto outperforms other non-IID FL methods, including the client clustering method HYPCLUSTER, adaptive federated learning (APFL), which interpolates between a global and local models [19], Loopless Local SGD (L2SGD), which applies global and local model average regularization [20], and MOCHA [6], which fits task-specific models through a multi-task objective. We refer readers to [8] for a comprehensive study of existing personalized FL methods.
@@ 2.2 Mode Connectivity and Solution Simplex
within which any point has a small loss, can be trained from randomly initialized endpoints. Here, $\{\boldsymbol{\theta}_{m}\in\mathbb{R}^{D}\}_{m=1}^{M+1}$ are the endpoints or vertices of the simplex, and $\Delta^{M}=\{\boldsymbol{\alpha}\in[0,1]^{M+1};\|\boldsymbol{\alpha}\|_{1}=1\}$ denotes the $M$-dimensional standard simplex. This simplex learning is performed by finding the endpoints that (approximately) minimize
@@
where $\mathcal{U}_{\mathcal{W}}$ denotes the uniform distribution on a set $\mathcal{W}$.
@@ 3.1 Federated Learning over Connected Modes (Floco)
The subregion assignments need to reflect the similarity between the clients. To this end, Floco expresses each client as a point in the standard simplex, based on the gradient update signals. Specifically, it applies the Euclidean projection onto the positive simplex [25] with the Riesz s-Energy regularization [26], which gives well spreaded projections that preserve the similarity between the client’s gradient signals as much as possible. Once the clients are projected onto the standard simplex as $\{\boldsymbol{\alpha}_{k}\in\Delta^{M}\}_{k=1}^{K}$, we assign the L1-ball with radius $r$ around $\boldsymbol{\alpha}_{k}$, i.e., $\mathcal{R}_{k}=\{\boldsymbol{\alpha}\in\Delta^{M};\|\boldsymbol{\alpha}-\boldsymbol{\alpha}_{k}\|_{1}\leq\rho\}$, to the $k$-th client. Note that the gradient update signals are informative for the subregion assignment only after the (global) model is trained to some extent. Therefore, the subregion assignment is performed after $\tau$ FL rounds are performed. Before the assignment, i.e., $t\leq\tau$, all clients train the whole standard simplex $\mathcal{R}_{k}=\Delta^{M},\forall k$, which corresponds to a simplex learning version of FedAvg.
@@
Although the simplex learning can be applied to all weight parameters, our preliminary experiment showed that simplex learning only of the weights in the last fully-connected layer (and the other weights are point-estimated) is sufficient. Therefore, our Floco only applies the simplex learning to the last layer, which gives other benefits including applicability to finetuning of pre-trained models, and significant reduction of computational and communication costs, see Section 4.3.
@@ 3.2 Client Gradient Projection onto Standard Simplex
Here we explain how to obtain the representations $\{\boldsymbol{\alpha}_{k}\in\Delta^{M}\}$ of the clients in the standard simplex such that similar clients are located close to each other, while all clients are well-spread across the simplex.
@@
The objective function in Eq.(8) is the Riesz s-Energy [26], a generalization of potential energy of multiple particles in a physical space, and therefore its minimizer correponds to the state where particles are well spread across the space. The minimization in the constraint (9) corresponds to the Euclidean projection onto the positive simplex , which forces $\{\boldsymbol{\beta}_{k}\}$ to keep the locations of the PCA projections $\{\boldsymbol{\kappa}_{k}\}$ of the clients. Fortunately, this minimization problem (for a fixed $z$) is convex, and can be efficiently solved (see Appendix A). We solve the main problem (8) by computing $\widehat{\boldsymbol{\beta}}_{k}(z)$ on a 1D grid in $z\in[0,1]$ with the interval $0.001$, and set the representations of the clients to $\boldsymbol{\alpha}_{k}=\frac{\widehat{\boldsymbol{\beta}}_{k}(\widehat{z})}{\widehat{z}}$, where $\widehat{z}$ is the minimizer of Eq.(8).
@@ 3.3 Communication Round: Local and Global Updates
In the $t$-th communication round, the server sends the current endpoints $\{\boldsymbol{\theta}_{m}^{t}\}_{m=1}^{M+1}$ to the participating clients $\mathcal{S}^{t}$. Then, each client $k\in\mathcal{S}^{t}$ draws one sample per mini-batch from the uniform distribution $\mathcal{A}=\{\boldsymbol{\alpha}_{b}\}_{b=1}^{B}\sim\mathcal{U}_{\mathcal{R}_{k}}$ on the assigned subregion,
and applies $T^{\prime}$ local updates,
@@
to the endpoints with $\boldsymbol{\alpha}$ sequentially chosen from $\mathcal{A}$111 Note that we do not rely on any regularizer that forces the diversity of the endpoints, as in [24], because, in Floco, the diversity of local client distributions prevents the simplex endpoints from collapsing to a single point. Here $\breve{\boldsymbol{\theta}}_{m}^{0}=\boldsymbol{\theta}_{m}^{t},\breve{\boldsymbol{\theta}}^{T^{\prime}}_{m}=\boldsymbol{\theta}^{t+1}_{m,k}$. The local updates $\{\Delta\boldsymbol{\theta}_{m,k}^{t+1}=\boldsymbol{\theta}^{t+1}_{m,k}-\boldsymbol{\theta}_{m}^{t}\}_{m=1}^{M+1}$ are sent back to the server, which updates the endpoints as
@@ 3.4 Inference
With the trained endpoints $\{\widehat{\boldsymbol{\theta}}_{m}=\boldsymbol{\theta}_{m}^{T}\}_{m=1}^{M+1}$, we simply use $\boldsymbol{w}_{\widehat{\boldsymbol{\alpha}}_{0}}(\{\widehat{\boldsymbol{\theta}}_{m}\}_{m=1}^{M+1})$ as the global model, where $\widehat{\boldsymbol{\alpha}}_{0}=\frac{1}{M+1}\boldsymbol{1}_{M+1}$ with $\boldsymbol{1}_{D}$ denoting the $D$-dimensional all one vector. For local models, we use $\{\boldsymbol{w}_{\widehat{\boldsymbol{\alpha}}_{k}}(\{\widehat{\boldsymbol{\theta}}_{m}\}_{m=1}^{M+1})\}_{k=1}^{K}$ where $\widehat{\boldsymbol{\alpha}}_{k}=\boldsymbol{\alpha}_{k}$.
@@ 4 Experiments
In this section, we experimentally show the advantages of Floco over the baseline methods.
@@ 4.1 Experimental Setting
Datasets and models
To evaluate Floco, we perform image classification on the CIFAR-10 [27] and the FEMNIST [28] datasets. For CIFAR-10, we train a simple CNN [29] from scratch, as well as fine-tune a ResNet-18 [30] pre-trained on ImageNet [31], as proposed in [32]. For FEMNIST, we train a simple CNN [29] from scratch as proposed in [29] as well as fine-tine a SqueezeNet [33] pre-trained on ImageNet as in [32]. We provide a table with the training hyperparemeters that we use for each data set/model setting in Appendix B.
Data heterogeneity for non-FL benchmarks
The FEMNIST dataset is an FL benchmark based on real data, where client heterogeneity is inherently embedded in the dataset. For CIFARO-10, we simulate statistical heterogeneity by two partitioning procedures. The first procedure by [34] partitions clients in equally sized groups and assigns each group a set of primary classes. Every client gets $z$ of its data from its group’s primary classes and $(100-z)$ from the remaining classes. We apply this method with $x=80$ for five groups and refer to this split as 5-Fold. For example, in CIFAR-10 5-Fold, 20 % of the clients get assigned 80 % samples from classes 1-2 and 20 % from classes 3-10. The second procedure, inspired by [35] and [36], draws the multinomial parameters of the client distributions $p_{k}(y)=\mathrm{Multi}(y;\boldsymbol{\phi}_{k})$ from Dirichlet, i.e., $\boldsymbol{\phi}_{k}\sim\mathrm{Dir}_{L}(\beta)$, where $\beta$ is the concentration parameter controlling the sparsity and heterogeneity—$\beta\to\infty$ concentrates the mass to the uniform distribution (and thus homogeneous), while small $10<\beta<$ 1 generates sparse and heterogeneous non-IID client distributions.
Baseline methods
We choose FedAvg [1], FedProx [13], and Ditto [14] as the baseline methods. The last two are the state-of-the-art for global FL and personalized FL, respectively.
Floco Hyperparameters
For simple CNN on the simulated non-IID splits Dir(0.5) and Five-Fold, we set $\tau=100,M=6,\rho=0.1$. For simple CNN on FEMNIST we set $\tau=200,M=3,\rho=0.3$. For pre-trained ResNet-18 on the simulated non-IID splits Dir(0.5) and Five-Fold we set $\tau=50,M=6,\rho=0.1$ and for the pre-trained SqueezeNet on FEMNIST we set $\tau=20,M=3,\rho=0.05$. We found those settings work well in our preliminary experiments, and conducted ablation study with other parameter settings in Appendix D. For the baselines, we follow the recommended parameter settings by the authors, which are detailed in Appendix B.
Evaluation criteria

### v4

@@ Abstract
Statistical heterogeneity in federated learning poses two major challenges: slow global training due to conflicting gradient signals, and the need of personalization for local distributions. In this work, we tackle both challenges by leveraging recent advances in linear mode connectivity — identifying a linearly connected low-loss region in the parameter space of neural networks, which we call solution simplex. We propose federated learning over connected modes (Floco), where clients are assigned local subregions in this simplex based on their gradient signals, and together learn the shared global solution simplex. This allows personalization of the client models to fit their local distributions within the degrees of freedom in the solution simplex and homogenizes the update signals for the global simplex training. Our experiments show that Floco accelerates the global training process, and significantly improves the local accuracy with minimal computational overhead in cross-silo federated learning settings.
@@ 1 Introduction
Federated learning (FL) [1] is a decentralized machine learning paradigm that facilitates collaborative model training across distributed devices while preserving data privacy. However, in typical real applications, statistical heterogeneity—non-identically and independently distributed (non-IID) data distributions at clients—makes it difficult to train well-performing models. To tackle this difficulty, various methods have been proposed, e.g., personalized FL [2], clustered FL [3], advanced client selection strategies [4], robust aggregation [5], and federated meta- and multi-task learning approaches [6]. These methods aim either at training a global model that performs well on the global distribution [7], or, as it is common in personalized FL, at training multiple client-dependent models each of which performs well on its local distribution [8]. These two aims often pose a trade-off—a model that shows better local performance tends to suffer from worse global performance, and vice versa. In this work, we aim to develop a FL method that improves local performance compared to state-of-the art methods without sacrificing global performance.
@@
Our approach leverages recent findings on mode connectivity [9, 10, 11]—the existence of low-loss paths in the parameter space between independently trained neural networks—and its applications [12]. These works show that minima for the same task are typically connected by simple low-loss curves, and that this connectivity benefits training for multi-task and continual learning. In particular, the authors show that embracing mode connectivity between models improves accuracy on each task and remedies the risk of catastrophic forgetting.
@@
Figure 1: Floco expresses each client as a point ($\star$ in the top-center plot) by projecting the gradient signals onto the simplex, so that similar clients are close to each other. In each communication round, each client uniformly samples points in the neighborhood of their projected point (top-right plot), and jointly train the solution simplex. The lower row shows the resulting test loss on the solution simplex, where the loss for the global distribution (left) is uniformly small, while the losses for individual local distributions (center for client 1 and right for client 2) are small around their projected points.
@@
In this paper, we leverage such effects, and propose federated learning over connected modes (Floco), where the clients share and together train a solution simplex—a linearly connected low-loss region in the parameter space. Specifically, Floco represents clients as points within the standard simplex based on the similarity between their gradients, and assigns each client a specific subregion of the simplex. Clients then participate in FL by sampling different models within their assigned subregions and sending back the gradient information to update the vertices of the global solution simplex (see Fig.1). This method facilitates collaborative training through the common solution simplex, while allowing for client-specific personalization according to their local data distributions.
@@
Our experiments show that Floco outperforms common FL baselines (FedAvg [1], FedProx [13]) and state-of-the-art personalized FL approaches (FedRoD [14], APFL [15], Ditto [16], FedPer [17]) on both local and global test metrics—without introducing significant computational overhead—in cross-silo FL settings. We also demonstrate additional benefits of Floco, including better uncertainty estimation, improved worst client performance, and smaller divergence of gradient signals.
@@
  • We propose Floco, a novel FL method that trains a solution simplex for mitigating the statistical heterogeneity of clients, and demonstrate its state-of-the-art performance for local personalized FL.
  • We propose a simple projection method to express clients as points in the standard simplex based on the gradient signals, and establish a procedure of subregion assignments.
  • We conduct experimental evaluations on semi-artificial and real-world FL benchmarks with detailed analyses of the behavior of Floco, which give insights into how the mechanism improves performance compared to the baselines.
We provide implementations of Floco in the FL frameworks FL-bench [18] and Flower [19]. Our code is publicly available: https://github.com/dennis-grinwald/floco.
@@ 2.1 Federated Learning
Here, $p(k)$ is the normalized population of data samples for the $k$-th client, $p_{k}(\boldsymbol{x},y)$ is the data distribution for the client $k$, and $f(g,(\boldsymbol{x},y))$ is the loss, e.g., cross-entropy, of the model $g$ on a sample $(\boldsymbol{x},y)\in\mathbb{R}^{I}\times\{1,\ldots,L\}$, where $I$ is the dimension of an input data sample. Global [20] and personalized [8] FL aim to approximate $g_{0}^{*}$ and $\{g_{k}^{*}\}_{k=1}^{K}$, respectively, by using the training data $\mathcal{D}=\{\mathcal{D}_{k}\}_{k=1}^{K}$ observed by the clients. Throughout the paper, we assume that all models are neural networks (NNs) $\widehat{y}=g_{k}(\boldsymbol{x};\boldsymbol{w}_{k})$ with the same architecture, and represent the model $g_{k}$ with its NN parameters $\boldsymbol{w}_{k}\in\mathbb{R}^{D}$, i.e., we hereafter represent $g_{k}(\boldsymbol{x};\boldsymbol{w}_{k})$ by $\boldsymbol{w}_{k}$ and thus denote, e.g., $F_{k}^{*}(g_{k})$ by $F_{k}^{*}(\boldsymbol{w}_{k})$. Let $N=\sum_{k=1}^{K}N_{k}$ be the total number of samples, where $N_{k}=|\mathcal{D}_{k}|$.
@@
where $\lambda$ controls the divergence from the global model. Ditto has been shown to outperform many other non-IID FL methods, including the client clustering method HYPCLUSTER, adaptive federated learning (APFL), which interpolates between a global and local models [25], Loopless Local SGD (L2SGD), which applies global and local model average regularization [26], and MOCHA [6], which fits task-specific models through a multi-task objective.
@@ 2.2 Mode Connectivity and Solution Simplex
within which any point has a small loss, can be trained from randomly initialized endpoints.
Here, $\{\boldsymbol{\theta}_{m}\in\mathbb{R}^{D}\}_{m=1}^{M+1}$ are the endpoints or vertices of the simplex, and $\Delta^{M}=\{\boldsymbol{\alpha}\in[0,1]^{M+1};\|\boldsymbol{\alpha}\|_{1}=1\}$ denotes the $M$-dimensional standard simplex. This simplex learning is performed by finding the endpoints that (approximately) minimize
@@
where $\mathcal{U}_{\mathcal{W}}$ denotes the uniform distribution on a set $\mathcal{W}$. During training, one model realization $w_{\alpha}$ from the simplex gets sampled and its gradient update wrt. the loss, e.g. cross-entropy, gets backpropagated to the simplex endpoints $\{\boldsymbol{\theta}_{m}\}_{m=1}^{M+1}$.
@@ 3.1 Federated Learning over Connected Modes (Floco)
The subregion assignments need to reflect the similarity between the clients. To this end, Floco expresses each client as a point in the standard simplex, based on the gradient update signals. Specifically, it applies the Euclidean projection onto the positive simplex [31] with the Riesz s-Energy regularization [32], which gives well spreaded projections that preserve the similarity between the client’s gradient signals as much as possible. Once the clients are projected onto the standard simplex as $\{\boldsymbol{\alpha}_{k}\in\Delta^{M}\}_{k=1}^{K}$, we assign the L1-ball with radius $\rho$ around $\boldsymbol{\alpha}_{k}$, i.e., $\mathcal{R}_{k}=\{\boldsymbol{\alpha}\in\Delta^{M};\|\boldsymbol{\alpha}-\boldsymbol{\alpha}_{k}\|_{1}\leq\rho\}$, to the $k$-th client. Note that the gradient update signals are informative for the subregion assignment only after the (global) model is trained to some extent. Therefore, the subregion assignment is performed after $\tau$ FL rounds are performed. Before the assignment, i.e., $t\leq\tau$, all clients train the whole standard simplex $\mathcal{R}_{k}=\Delta^{M},\forall k$, which corresponds to a simplex learning version of FedAvg.
@@
Although the simplex learning can be applied to all parameters, our preliminary experiment showed that applying simplex learning only of the parameters in the last fully-connected layer (while point-estimating the other parameters) is sufficient. Therefore, our Floco only applies the simplex learning to the last layer, which gives other benefits including applicability to fine-tuning of pre-trained models, and significant reduction of computational and communication costs, as shown in Section 4.3.
@@ 3.2 Client Gradient Projection onto Standard Simplex
We explain how to obtain the representations $\{\boldsymbol{\alpha}_{k}\in\Delta^{M}\}$ of the clients in the standard simplex such that similar clients are located close to each other, while all clients are well-spread across the simplex.
@@
The objective function in Eq. (8) is the Riesz s-Energy [32], a generalization of potential energy of multiple particles in a physical space, and therefore its minimizer correponds to the state where particles are well spread across the space. The minimization in the constraint (9) corresponds to the Euclidean projection onto the positive simplex [31], which forces $\{\boldsymbol{\beta}_{k}\}$ to keep the locations of the PCA projections $\{\boldsymbol{\kappa}_{k}\}$ of the clients. Fortunately, this minimization problem (for a fixed $z$) is convex, and can be efficiently solved (see Appendix A). We solve the main problem (8) by computing $\widehat{\boldsymbol{\beta}}_{k}(z)$ on a 1D grid in $z\in[0,1]$ with the interval $0.001$, and set the representations of the clients to $\boldsymbol{\alpha}_{k}=\frac{\widehat{\boldsymbol{\beta}}_{k}(\widehat{z})}{\widehat{z}}$, where $\widehat{z}$ is the minimizer of Eq. (8).
@@ 3.3 Communication Round: Local and Global Updates
In the $t$-th communication round, the server sends the current endpoints $\{\boldsymbol{\theta}_{m}^{t}\}_{m=1}^{M+1}$ to the participating clients $\mathcal{S}^{t}$. Then, each client $k\in\mathcal{S}^{t}$ draws one sample per mini-batch from the uniform distribution $\mathcal{A}=\{\boldsymbol{\alpha}_{b}\}_{b=1}^{B}\sim\mathcal{U}_{\mathcal{R}_{k}}$ on the assigned subregion and applies $T^{\prime}$ local updates,
@@
to the endpoints with $\boldsymbol{\alpha}$ sequentially chosen from $\mathcal{A}$111 Note, that we do not rely on any regularizer that forces the diversity of the endpoints, as in [30]. In Floco, the diversity of local client distributions prevents the simplex endpoints from collapsing to a single point. Here $\breve{\boldsymbol{\theta}}_{m}^{0}=\boldsymbol{\theta}_{m}^{t},\breve{\boldsymbol{\theta}}^{T^{\prime}}_{m}=\boldsymbol{\theta}^{t+1}_{m,k}$. The local updates $\{\Delta\boldsymbol{\theta}_{m,k}^{t+1}=\boldsymbol{\theta}^{t+1}_{m,k}-\boldsymbol{\theta}_{m}^{t}\}_{m=1}^{M+1}$ are sent back to the server, which updates the endpoints as
@@
3.4 Floco+
We can further enhance the personalized FL performance of Floco by additionally fine-tuning a local model as in Ditto [16]. In this extension, called Floco+, each client personalizes the global endpoints $\{\widehat{\boldsymbol{\theta}}_{m}^{0}=\boldsymbol{\theta}_{m}\}_{m=1}^{M}$ by local gradient descent to minimize the Ditto objective, i.e.,
@@ 3.5 Inference
With the trained endpoints $\{\widehat{\boldsymbol{\theta}}_{m}=\boldsymbol{\theta}_{m}^{T}\}_{m=1}^{M+1}$, we simply use $\boldsymbol{w}_{\widehat{\boldsymbol{\alpha}}_{0}}(\{\widehat{\boldsymbol{\theta}}_{m}\}_{m=1}^{M+1})$ as the global model, where $\widehat{\boldsymbol{\alpha}}_{0}=\frac{1}{M+1}\boldsymbol{1}_{M+1}$ with $\boldsymbol{1}_{D}$ denoting the $D$-dimensional all one vector. For local models, we use $\{\boldsymbol{w}_{\widehat{\boldsymbol{\alpha}}_{k}}(\{\widehat{\boldsymbol{\theta}}_{m}\}_{m=1}^{M+1})\}_{k=1}^{K}$ where $\widehat{\boldsymbol{\alpha}}_{k}=\boldsymbol{\alpha}_{k}$. For Floco+, we fine-tune the corresponding subspace regions $\mathcal{R}_{z_{k}}$ for $E$ local epochs.
@@ 4 Experiments
In this section, we experimentally show the advantages of Floco and Floco+ over the baselines.
@@ 4.1 Experimental Setting
Datasets and models.
To evaluate our method, we perform image classification on the CIFAR-10 [33] and FEMNIST [34] datasets. For CIFAR-10, we train a CNN (CifarCNN) from scratch, following [35], and fine-tune a ResNet-18 [36] pre-trained on ImageNet [37], as in [38]. For FEMNIST, we train a CNN (FemnistCNN) from scratch, as in [1], and fine-tune a SqueezeNet [39] pre-trained on ImageNet, following [38]. We provide a table with the training hyperparameters that we use for each dataset/model setting in Appendix B.
Data heterogeneity for non-FL benchmarks.
The FEMNIST dataset is an FL benchmark based on real data, where client heterogeneity is inherently embedded in the dataset. For CIFAR-10, we simulate statistical heterogeneity by two partitioning procedures. The first procedure by [40] partitions clients in equally sized groups and assigns each group a set of primary classes. Every client gets $q$ of its data from its group’s primary classes and $(100-q)$ from the remaining classes. We apply this method with $q=80$ for five groups and refer to this split as 5-Fold. For example, in CIFAR-10 5-Fold, 20 % of the clients get assigned 80 % samples from classes 1-2 and 20 % from classes 3-10. The second procedure, inspired by [41] and [42], draws the multinomial parameters of the client distributions $p_{k}(y)=\mathrm{Multi}(y;\boldsymbol{\phi}_{k})$ from Dirichlet, i.e., $\boldsymbol{\phi}_{k}\sim\mathrm{Dir}_{L}(\beta)$, where $\beta$ is the concentration parameter controlling the sparsity and heterogeneity—$\beta\to\infty$ concentrates the mass to the uniform distribution (and thus homogeneous), while small $10<\beta<$ 1 generates sparse and heterogeneous non-IID client distributions.
Baseline methods.
Besides FedAvg [1] and FedProx [21] for global FL, we chose FedRoD [14], APFL [15], Ditto [16], and FedPer [17] as state-of-the-art personalized FL baselines.
Floco Hyperparameters.
For CifarCNN on the simulated non-IID splits Dir(0.3)/Five-Fold, we set $\tau=250,M=20/10,\rho=0.1$. For FemnistCNN on FEMNIST we set $\tau=250,M=10,\rho=0.5$. For pre-trained ResNet-18 on the simulated non-IID splits Dir(0.3)/Five-Fold we set $\tau=50,M=20/10,\rho=0.1$ and for the pre-trained SqueezeNet on FEMNIST we set $\tau=250,M=3,\rho=0.5$. We found those settings work well in our preliminary experiments, and conducted ablation study with other parameter settings in Appendix D. For the baselines, we follow the recommended parameter settings by the authors, which are detailed in Appendix B.
Evaluation criteria.